OpenStudy (anonymous):

what is cot x= 3 (square root) 3

OpenStudy (anonymous):

Cotx = 1/(tanx) So when cot=root3, tan= 1/root3 Remember the triangle with sides: 1, 2 and root3 Draw that out, with the angles, remembering that tan= opposite/adjacent. You'll find that the tan30=1/root3 So 1/(1/root3) = cotx Cotx = root3 Therefore, x=30